Example: For the multiple choice test example (10 questions, each with 8 choices), what is the probability that the person gets 2 or 3 questions correct?
Answer #1

X~ binomial (n,p)  

Where n = 10, p = 1/8

Binomial probability distribution is

P(X) = nCx px (1-p)n-x

So,

P(X = 2 or 3) = P( X = 2) + p( x = 3)

= 10C2 (1/8)2 (7/8)8+10C3 (1/8)3 (7/8)7

= 0.3336